# First Observation of Antiproton Annihilation At Rest on Argon in the LArIAT Experiment

### v1

@@ V Final-State Reconstruction Results
As with the incoming antiproton candidates, the particle tracks leaving of the annihilation vertex are passed through the PID algorithm to determine their particle type with the results for the data and simulated events shown in Fig. 9. The simulation was normalized to the number of data antiproton events to allow direct comparisons. Both the data and simulation effectively agree within statistical uncertainty. Half of the outgoing particles have an end point within 1 cm from an edge indicating that they are considered to be exiting the LArTPC. This makes it impossible to perform a definitive PID based on the d/d of these tracks and instead contribute to the last bin in Fig. 9.
The final-state outgoing track multiplicity is obtained in two different ways: by hand-counting the number of tracks leaving the annihilation vertex, and by counting the number of reconstructed tracks determined by the automated tracking algorithm. The first method is included due to the reconstruction failures mentioned in Sec. IV.1. This allows for a comparison between the two methods and is possible due to the small size of the final candidate sample in this analysis. Due to complications from manually clustering multiple tracks emanating from a common vertex, hits near the vertex were often excluded, as explained in Sec. IV.2. Therefore, a generous maximum track-to-vertex distance of 5 cm was allowed for outgoing track candidates, with a required span of at least several wires to ensure they extend far enough out from the vertex region.
The track multiplicity for the two methods for the simulation and data events is shown in Fig. 10. The calculated track multiplicity agrees well with the simulation within statistical uncertainty.
A summary of the computed means, uncertainty on the means, and the root mean squares (RMS) values for both cases is found in Table 2 where the mean for the reconstructed counting method is slightly lower than for the hand-scanned due to reconstruction failures and inefficiencies.
@@
Automated reconstruction |
Antiproton annihilation may also produce and [13, 14], both of which decay into photon showers (). They decay promptly which can result in a pair of showers pointing directly back to the annihilation vertex at some distance. Showers pointing back to the vertex are counted through hand-scanning but the ones pointing back at an outgoing track are not counted as they are considered to originate from one of the outgoing particle tracks rather than directly from the annihilation. Due to the conversion length of (greater than 29 cm [36]), and the small size of LArIAT, they can exit the LArTPC before producing a visible shower. The simulation predicts that 40% of the exit the LArTPC before interacting. This results in an observed lower limit on the number of showers coming from the annihilations. Similar to the track multiplicity, the shower multiplicity for data (simulation) corresponds to () showers with RMS of 1.8 (1.6) with the distributions shown in Fig. 11. Here as well, the shape of the distributions agree within statistical uncertainty.
@@ VI Conclusions
Using a LArTPC detector in a test beam, LArIAT has observed the first antiproton annihilation at rest interactions on argon . Using LArIAT’s Run II negative polarity data, a total of 11 candidates of antiproton annihilation at rest are identified using various beamline and LArTPC selections. Within our statistics, data and simulation distribution of the final-state multiplicities and particle type for surface antiproton annihilation at rest using the FTF model as implemented in Geant4 agree well. For data, a mean of tracks per annihilation is measured with RMS of 1.3. Similarly, measurements of shower multiplicity also agree between data and MC, with data showing a mean of showers within LArIAT’s active volume with RMS of 1.8. These results strengthen the confidence in the robustness and reliability of current model-building efforts for - annihilation inside the nucleus by adding a new nucleus to the list of observed and measured antiproton annihilation at rest events.
VII Acknowledgements
@@ Appendix A Selected Data Antiproton Candidates
Figure 12 show a selection of representative data and MC event displays for antiproton annihilation at rest. The highly ionizing track traveling from the middle-left towards the center of the panels are the incoming antiproton marked in white with and arrow for the direction. At the end of the antiproton’s Bragg peak, outgoing tracks are seen exiting directly from the annihilation vertex, while outgoing showers exhibit a gap between the vertex and their starting point but are still oriented directly towards the vertex. The number of hand-scanned tracks (Ntracks) and showers (Nshowers) for these events are indicated in the event displays. In the case of (a) and (c), the relatively straight tracks near the top half of the collection and induction plane panes are high-energy ‘halo’ muons produced from the pion decays in the secondary beam far upstream of the LArIAT tertiary beamline, which can contaminate beamline-triggered readouts with multiple extra tracks unassociated with the tertiary beamline.

### v2

@@ V Final-State Reconstruction Results
As with the incoming antiproton candidates, the particle tracks leaving of the annihilation vertex are passed through the PID algorithm to determine their particle type with the results for the data and simulated events shown in Fig. 8. The simulation was normalized to the number of data antiproton events to allow direct comparisons. The data and simulation are consistent within statistical uncertainty. Half of the outgoing particles have an end point within 1 cm from an edge in which case they are considered to be exiting the LArTPC. This makes it impossible to perform a definitive PID based on the d/d of these tracks and these contribute to the last bin in Fig. 8.
The final-state outgoing track multiplicity is obtained in two different ways: by hand-counting the number of tracks leaving the annihilation vertex, and by counting the number of reconstructed tracks determined by the semi-automated algorithm. This allows for a comparison between the two methods and is possible because of the small size of the final candidate sample in this analysis. Due to complications from manually clustering multiple tracks emanating from a common vertex, hits near the vertex were often excluded, as explained in Sec. IV.2. Therefore, a generous maximum track-to-vertex distance of 2 cm was allowed for outgoing track candidates, with a required span of five wires to ensure that tracks extend far enough from the vertex region.
The track multiplicity distributions comparing simulation to data for both the hand-scanning and semi-automated reconstruction methods are shown in Fig. 9. The data distributions are consistent with the simulation within statistical uncertainty.
A summary of the calculated means, the uncertainty of the means and the root mean squares (RMS) values for both track multiplicities for the hand-scanning and semi-automated reconstruction is shown in Table 2. The mean for the semi-automated reconstructed counting method is slightly lower than for the hand-scanned due to reconstruction failures and inefficiencies. All calculated means agree within uncertainties.
@@
| Semi-automated reconstruction |
Antiproton annihilation may also produce and [13, 14], both of which decay into photon showers (). These particles decay promptly which can result in a pair of showers pointing directly back to the annihilation vertex at some distance. Showers pointing back to the vertex are counted through hand-scanning but the ones pointing back to an outgoing track are not counted as they are considered to originate from one of the outgoing particle tracks rather than directly from the annihilation. Due to the conversion length of the (greater than 29 cm [39]), and the small size of LArIAT, can exit the LArTPC before producing a visible shower. The simulation predicts that 40% of the exit the LArTPC before interacting. This results in an observed lower limit on the number of showers coming from the annihilations. Similar to the track multiplicity, the mean shower multiplicity for data (simulation) corresponds to () showers with RMS of 1.8 (1.6). The distributions are shown in Fig. 10 and data and simulation are consistent within statistical uncertainty.
@@ VI Conclusions
LArIAT has observed the first antiproton annihilation at rest interactions on argon using a LArTPC detector in a test beam. A total of 11 candidates of antiproton annihilation at rest are identified in LArIAT’s Run II negative polarity data using various beamline and LArTPC selections. Distributions of final state particle multiplicity and type were compared to MC simulations using Geant4’s FTF model for surface nucleon annihilation. For data, a mean of tracks per annihilation is measured with a RMS of 1.3 for the hand-scanned method and tracks with a RMS of 1.2 for the semi-automated method. Measurements of shower multiplicity are also consistent between data and simulation, with data showing a mean of showers within LArIAT’s active volume with RMS of 1.8. The predictions from the models used in Geant4 demonstrate consistency with the data, as they fall within the statistical uncertainty for the low-statistics sample used in this paper. The methods outlined in this paper offer a foundation for higher-statistics measurements in larger LArTPCs, which could further assess the models’ agreement and potentially inform improvements in modeling approaches by adding a new nucleus to the list of observed antiproton annihilation at rest events.
VII Acknowledgments
@@ Appendix A Selected Data Antiproton Candidates
Figure 11 shows a selection of representative data and MC event displays for antiproton annihilation at rest. The highly ionizing track traveling from the middle-left towards the center of the panels are the incoming antiproton marked in white with and arrow for the direction. At the end of the antiproton’s Bragg peak, outgoing tracks are seen exiting directly from the annihilation vertex, while outgoing showers exhibit a gap between the vertex and their starting point but are still oriented directly towards the vertex. The number of hand-scanned tracks (Ntracks) and showers (Nshowers) for these events are indicated in the event displays. In the case of (a) and (c), the relatively straight tracks near the top half of the collection and induction plane panes are high-energy muons produced from the pion decays in the secondary beam far upstream of the LArIAT tertiary beamline, which can contaminate beamline-triggered readouts with multiple extra tracks unassociated with the tertiary beamline.